OpenStudy (marie0424):

Grant's credit card has an APR of 11.28%, and it just changed its compounding period from monthly to daily. What will happen to the effective interest rate charged to Grant?

jimthompson5910 (jim_thompson5910):

The formula you'll use is found on this page https://en.wikipedia.org/wiki/Effective_interest_rate The formula is \[\Large r = \left(1+\frac{i}{n}\right)^n-1\]

jimthompson5910 (jim_thompson5910):

r = effective annual rate i = nominal rate n = compounding frequency

jimthompson5910 (jim_thompson5910):

`Grant's credit card has an APR of 11.28%` so that means i = 11.28/100 = 0.1128 r = unknown which is what you want to find out n = 12 for the monthly compounding n = 365 for the daily compounding

jimthompson5910 (jim_thompson5910):

let me know if that helps you get started